OpenStudy (anonymous):

A spotlight on the ground shines on a wall 12 m away. If a man 2 m tall walks from the spotlight toward the building at a speed of 2.1 m/s, how fast is the length of his shadow on the building decreasing when he is 4 m from the building?

OpenStudy (anonymous):

I made a diagram

OpenStudy (anonymous):



OpenStudy (anonymous):

I can't get the right answer however.

OpenStudy (anonymous):

I also know \[\frac{ dx }{ dt }=2.1m/s\] and I have to use the concept of similar triangles.

OpenStudy (phi):

I would let x be the distance from the spotlight (that way dx/dt has the correct sign, meaning x gets bigger with increasing time) You called the height of the shadow y. I would move the man closer to the wall (dx amount) and note that the shadow gets shorter. So expect dy/dx to be negative. The question is asking for dy/dt, and they give you dx/dt. This suggests using \[ \frac{dy}{dt}= \frac{dx}{dt}\cdot \frac{dy}{dx} \] We need to find dy/dx. That means we need an equation the relates x and y. Similar triangles will work.  we see xy= 24 differentiate with respect to x and solve for dy/dx Can you finish?